### 18 Mar 2024 v1

# Exploring Facial Expression Recognition through Semi-Supervised Pretraining and Temporal Modeling

Abstract
Facial Expression Recognition (FER) plays a crucial role in computer vision and finds extensive applications across various fields. This paper aims to present our approach for the upcoming 6th Affective Behavior Analysis in-the-Wild (ABAW) competition, scheduled to be held at CVPR2024.. In the facial expression recognition task, The limited size of the FER dataset poses a challenge to the expression recognition model’s generalization ability, resulting in subpar recognition performance. To address this problem, we employ a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled face data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to address the problem of category imbalance in the dataset and the possible data bias in semi-supervised learning. Moreover, , to further compensate for the limitation and bias of features obtained only from static images, we introduced a Temporal Encoder to learn and capture temporal relationships between neighboring expression image features. In the 6th ABAW competition, our method achieved outstanding results on the official validation set, a result that fully confirms the effectiveness and competitiveness of our proposed method.
1 Introduction
The goal of Facial Expression Recognition (FER) is to identify the emotional state of an individual by analyzing facial images or videos. It is a broad research field spanning multiple fields such as machine learning, image processing, psychology, etc., with a wide range of applications, including safe driving, intelligent monitoring, and human-computer interaction. In view of its diverse applications, it is extremely important to establish a robust FER system. The expression recognition task is a classic problem in the field of pattern recognition, which usually involves the classification of six basic emotions: happiness, surprise, sadness, anger, disgust and fear. The facial expression recognition community has made tremendous progress in recent years. State-of-the-art FER methods achieve very good results on numerous public datasets, such as RAF-DB[23], SFEW[34] and AffectNet[24].
To stimulate interdisciplinary collaboration and address pivotal research inquiries spanning affective computing, machine learning, and multi-modal signal processing, Kollias et al. have spearheaded the Affective Behavior Analysis in-the-wild (ABAW) initiative. The 6th iteration of this workshop and competition is slated to align with the IEEE CVPR conference in 2024[35, 16, 12, 17, 11, 18, 19, 13, 9, 10, 14, 15].
Traditional fully supervised Facial Expression Recognition (FER) methods rely on the availability of large volumes of high-quality labeled data to fine-tune model precision [28, 3, 2]. However, mainstream training datasets often suffer from class imbalance. Most fully supervised models tend to accurately identify majority classes, thus reducing the model’s accuracy for minority classes. Sometimes minority classes constitute less than 10% of the data, and this disparity in data volume makes it challenging for models to learn fairly across all categories. Financial and logistical constraints in acquiring extensive labeled FER data hinder the expansion of training repositories. In contrast, the volume of data for Face Recognition (FR) surpasses that of FER. Making a leap to augment samples from FR data to aid models in learning the FER task could be highly beneficial. Addressing the class imbalance issue and the discrepancy in data distribution between FR and FER data to effectively utilize FR data in support of FER models presents an urgent challenge.
In this study, we propose a two-phase methodology to enhance the recognition and analysis of facial expressions. The first phase, known as the spatial pre-training phase, plays a crucial role in preparing the model for subsequent tasks. During this phase, we leverage the power of semi-supervised learning techniques to generate pseudo-labels for expression categories using unlabeled face data. This approach ensures a sufficiently large training corpus, allowing the model to effectively extract robust facial expression features. To tackle the challenges of category imbalance in the dataset and potential data bias during semi-supervised learning, we adopt two strategies. First, we uniformly sample labeled facial expression instances to address the category imbalance. Additionally, we employ a debiased feedback learning strategy to mitigate the impact of potential data bias. These strategies collectively contribute to training a more robust facial expression recognizer. Moving to the second phase, the temporal learning phase, we aim to further improve the recognition and analysis of facial expressions by capturing the temporal dynamics. In this phase, we freeze the facial expression recognizer trained in the first phase, which has already acquired strong spatial representation capabilities. To incorporate temporal information, we introduce a temporal encoder that learns the temporal relationships between neighboring expression image features. By considering the temporal aspect, we compensate for the inherent feature bias obtained solely from static images. This integration of temporal dynamics enables more accurate and comprehensive dynamic recognition and analysis of facial expressions.
To sum up, our contributions can be summarized as:
  • To address the problem of scarcity of facial expression data, we applied a semi-supervised learning technique to generate expression category pseudo-labels for unlabeled face data. At the same time, we uniformly sampled the labeled facial expression samples and implemented a debiased feedback learning strategy to solve the problem of category imbalance in the dataset and the possible data bias in the semi-supervised learning, which improves the robustness and generalization ability of the model.
  • In order to compensate for the limitations and biases of features acquired only from static images, we introduce a temporal encoder to learn and capture temporal relationships between neighboring expression image features. This strategy aims to enhance the model’s ability to recognize and analyze the dynamic changes in facial expressions, and achieve more accurate dynamic facial expression recognition.
  • In the 6th ABAW competition, our algorithm achieved outstanding results on the official validation set, a result that fully confirms the effectiveness and competitiveness of our proposed method.
Figure 1: Framework Description. Our approach is mainly divided into a spatial phase and a temporal phase. (1). The goal of the spatial phase is to expand face expression data by mining large-scale unlabeled faces through a semi-supervised algorithm. (2). The goal of the temporal phase is to do temporal feature enhancement of the image features extracted by the student network in the first phase by means of a temporal encoder, so as to improve the accuracy of recognizing the dynamic facial expressions in the video.
2 Related Work
2.1 Facial Expression Recognition
The facial expression recognition task is a classic task in the field of pattern recognition, and the approach[29, 26, 39, 3, 2, 33] of making full use of fully supervised data once gained very great progress in the field of FER. In recent years, attention has been focused on extending the dataset to obtain larger scale datasets. To solve the label confusion problem between different expression recognition datasets. IPA2LT [37] is the first work to address the annotation inconsistency in different facial expression datasets. They proposed LTNet embedded with a scheme of discovering the latent truth from multiple inconsistent labels and the input images. Ada-CM [22] is the first solution to explore the dynamic confidence margin in Semi-Supervised Deep Facial Expression Recognition. They designed an adaptive confidence margin to dynamically learn on all unlabeled data for the model’s training and conducted a feature-level contrastive objective to learn effective features by applying the InfoNCE[30] loss. Face2Exp [36] proposed the Meta-Face2Exp framework to extract de-biased knowledge from auxiliary FR data through the meta optimization framework. utilized unlabeled face data to enhance expression recognition through the meta optimization framework.
2.2 Learning with Unlabeled Data
An important direction for learning methods using unlabeled data is semi-supervised learning. A popular class of Semi-Supervised Learning methods is to generate an artificial label for an unlabeled image and train the model to predict that artificial label on the input of an unlabeled image [21, 31]. Similarly, consistent regularization [1, 20, 25] uses the predictive distribution of the model to obtain an artificial label after randomly modifying the input or model function. An artificial label is obtained using the predictive distribution of the model after randomly modifying the input or the model function. Fixmatch[27] combines the advantages of these two methods, which use weak data augmentation and strong data augmentation of samples to achieve consistent regularization and obtain pseudo-labeled data by samples with confidence levels greater than a threshold. But the problem of this approach is that his modeling ability is limited in the early stage of model training because the threshold value is fixed. To solve this problem, Flexmatch[38] proposed Curriculum Pseudo Labeling (CPL), a curriculum learning approach to leverage unlabeled data according to the model’s learning status. It can flexibly adjust thresholds for different classes at each time step to let pass informative unlabeled data and their pseudo labels. In addition, Dash[32] performs selection by retaining only samples with losses less than a given threshold in each update iteration, which is dynamically adjusted through iterations.
3 Method
3.1 Overview
As shown in Figure.1, our approach uses multiple techniques to improve facial expression recognition (FER) in two phases. The first phase, the spatial phase, uses semi-supervised learning to increase the FER dataset by using facial recognition (FR) data. This helps in training a better image feature extractor. We use two neural networks, a teacher and a student, with the same structure but different weights.In this phase, we improve the networks by using a debiasing method. This involves comparing biased FR data (unlabeled) with debiased FER data (labeled) to create pseudo-labels. The teacher network learns from balanced FER data and makes these pseudo-labels, which the student network then uses to learn from the FR data, adjusting based on feedback to improve accuracy.In the training, the teacher and student networks update alternately with the debiasing method to better process FR data, improving accuracy and debiasing. The second phase, the temporal phase, uses the student network (fixed after the first phase) to extract image features. To address biases from static images, we use a temporal encoder to understand time-related feature relationships, improving system accuracy. This enables dynamic recognition and analysis of facial expressions in videos, with a classifier providing frame-by-frame predictions.
3.2 Data Pre-process
3.2.1 Data Sampling
To achieve class balance, we meticulously sample the labeled Facial Expression Recognition (FER) data, ensuring an equal distribution of samples across each expression category. This strategic sampling method allows the model to learn features that are more evenly distributed among classes, significantly contributing to the de-biasing process in Facial Recognition (FR) data analysis.
3.2.2 Data Augmentation
To enhance the diversity and robustness of our dataset, we implement two distinct data augmentation strategies. In the case of Weakly-Augmented (WA) operations, our approach primarily encompasses horizontal flipping, color jittering, and similar techniques to introduce subtle variations. On the other hand, for Strongly-Augmented (SA) operations, we adopt RandAugment[4], a method known for its efficacy in generating more significant transformations while maintaining the integrity of the data. These augmentation strategies are meticulously chosen to ensure that, despite the alterations, the unlabeled samples retain their semantic consistency.
3.3 Semi-Supervised Training
The focus of this phase is to utilize semi-supervised learning to extend the FER dataset by using the Facial Recognition (FR) dataset. This expansion helps train an efficient static image feature extractor. In this phase, we employ two neural networks with identical structures but independent weights: the teacher network (T) and the student network (S).
To optimize network performance in this phase, we employ a debiasing mechanism as a core strategy. This involves analyzing the disparities between biased FR data (which is unlabeled) and debiased FER data (which is labeled) to generate pseudo-labels. By sampling FER data in a category-balanced manner, the teacher network learns and generates these pseudo-labels. Subsequently, the student network employs these pseudo-labels to train on the FR data, continuously adjusting itself based on feedback to enhance recognition accuracy.During the training process, the teacher network and the student network are alternately updated, coupled with the de-biasing strategy. This iterative updating aims to improve both the accuracy and de-biasing effect of processing FR data. Consequently, this enhances the prediction ability of the student network.
3.3.1 Student Network
To enhance FER, the student network leverages the rich and comprehensive diversity of large-scale unlabeled FR data. During training, the student network utilizes this unlabeled FR data along with pseudo-labels generated by the teacher’s network. The training process involves encouraging both networks to predict similar conditional classification distributions for the unlabeled FR data, achieved through the use of the $\mathcal{L}_{u}$ loss function. The expression for $\mathcal{L}_{u}$ is given as:
$\mathcal{L}_{u}=\mathrm{CE}(\hat{y}_{FR},\mathcal{S}(x_{FR};\theta_{\mathrm{s}})).$ (1)
where $\theta_{\mathrm{s}}$ is the parameter of the student network, and the $\mathrm{CE}$ denotes the cross-entropy loss.
As the student network undergoes updates, the pseudo-labeling dynamically changes throughout the training process. The parameters $\theta_{\mathrm{s}}$ of the student network are updated during the meta-training phase. In the meta-testing phase, the balanced FER dataset (which was used to train the teacher network) is employed to estimate the cognitive differences between the biased FR data and the de-biased FER.
3.3.2 Teacher Network
A sampling module, denoted as $\mathrm{Smp(\cdot)}$, is utilized to guarantee a balanced class distribution in the Facial Expression Recognition (FER) dataset. To achieve this, an equal number of samples from each facial expression category were randomly selected, thereby ensuring a balanced class representation for training the teacher network. The learning process integrates three specific types of loss functions: supervised loss, consistency loss, and feedback loss, which collectively provide effective guidance to the teacher network. This process is represented by the following equation:
$\mathcal{L}_{\mathcal{T}}=\mathcal{L}_{s}+\mathcal{L}_{c}+\mathcal{L}_{f}.$ (2)
Specifically, the supervision loss and consistency loss apply only to the teacher network, while the feedback loss takes into account the predictive distribution of the meta-test for the student network.
For the supervised learning, aimed at minimizing the supervised loss function on a balanced and labeled Facial Expression Recognition (FER) dataset, the supervised loss function $\mathcal{L}_{s}$ is formulated as follows:
$\mathcal{L}_{s}=\mathsf{C}\text{E}(y_{FER},\mathcal{T}(x_{FER};\theta_{\mathrm{t}})).$ (3)
where $\theta_{\mathrm{t}}$ is a parameter of the teacher network and $\mathsf{C}\text{E}$ denotes the cross-entropy loss.
For consistency learning,the teacher network requires that the original image and the augmented counterpart have close class-conditional distributions with a consistency loss function $\mathcal{L}_{c}$:
$\mathcal{L}_{c}=\mathrm{CE}(\mathcal{T}(x_{FR};\theta_{\mathrm{t}}),\mathcal{T}(\mathrm{Aug}(x_{FR});\theta_{\mathrm{t}})).$ (4)
The $\mathrm{Aug(\cdot)}$ denotes data augmentation, this approach utilizes intensive data augmentation techniques, including rotation, removal, and pixel-level image manipulation.
For the feedback learning, the process involves estimating feedback based on cognitive discrepancies between the FR (Facial Recognition) and FER (Facial Expression Recognition) datasets. This feedback is utilized to refine the parameters of the teacher network, with the feedback loss function $\mathcal{L}_{f}$ denoted by:
$\mathcal{L}_{f}=f\cdot\mathrm{CE}(\hat{y}_{FR},\mathcal{T}(x_{FR};\theta_{\mathrm{t}})).$ (5)
where the definition of the feedback coefficient $f$ can be expressed as:
$f=\eta_{\mathcal{S}}\cdot(\nabla_{\theta_{\mathcal{S}}^{(t+1)}}\mathrm{CE}(y_{FER},\mathcal{S}(x_{FER};\theta_{\mathcal{S}}^{(t+1)}))^{\top}\cdot$ (6)
$\nabla_{\theta_{\mathcal{S}}}\mathrm{CE}(\hat{y}_{FR},\mathcal{S}(x_{FR};\theta_{\mathcal{S}}^{(t)}))).$
The $f$ is denoted as the dot product of two terms. First term: the gradient of the new student network on the debiased FER data. Second term: gradient of the old student network on biased FR data
3.3.3 Debiasing Mechanism
The behavior of new adaptation network on balanced FER data is used as the measurement. Specifically, when adaptation network on FR data and new adaptation network on FER data have the same cognitive (i.e., sign of their gradients), we obtain positive feedback coeffi- cients (i.e., positive value for Lf ), which encourages the up- date of base network by using the current direction of gradients. When adaptation network on FR data and new adaptation network on FER data have different cognitive (i.e., sign of their gradients), feedback coefficients will give a negative sign, which punishes the update of base network by using the adverse direction of current gradients. In this way, feedback is used as the reward signal that goes back through the base network, determining how parameters of the base network impact the gradient of the adaptation network for debiased feature extraction. As a result, the learning of adaptation network on biased FR data will perform de-biased behavior that is consistent with the evaluation procedure on balanced FER data.
3.4 Temporal encoder
The Transformer-based Temporal Encoder is a novel neural network architecture that we have specifically developed to capture temporal relationships between image features, aiming to address feature biases obtained solely from static images. This design leverages the self-attention mechanism of the Transformer architecture to effectively process sequence data and learn temporal dependencies within image sequences. By applying self-attention in the temporal dimension, the Temporal Encoder enables the model to emphasize key frames and features in an image sequence, facilitating the extraction of spatio-temporal features.
3.5 Post-Process
Given that the Aff-Wild2 dataset is constructed from consecutive frames of videos, and considering that the formation of facial expressions unfolds over a period of time, it follows logically that significant alterations in facial expressions are unlikely to occur within a narrow sequence of adjacent frames. In light of this, we implemented a sliding window technique to refine the predictive outcomes, aiming to enhance the consistency of the expression labels. This process involves aggregating the frequencies of all predicted labels within each window. Subsequently, the label that predominates in frequency within a given window is designated as the representative expression for all frames encompassed by that window. By applying this sliding window approach across the entire dataset, we effectively facilitate the smoothing of predicted expression labels, thereby achieving a more coherent and accurate representation of facial expressions throughout the dataset.
Table 1: Ablation study results on the official validation set.
Method Aff-Wild2 AffectNet and ExpW MS1MV2 Temporal Encoder Post-Process F1 Score (%)
Baseline $\checkmark$ 27.65
SSL $\checkmark$ $\checkmark$ 40.96
SSL $\checkmark$ $\checkmark$ 41.57
SSL + Temporal $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 43.77
SSL + Temporal+ Post-process $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 45.43
4 Experiment
In this section, we will provide a detailed description of the used datasets, the experiment setup, and the experimental results.
4.1 Datasets
FER Datasets. The 6th Workshop and Competition on Affective Behavior Analysis in-the-wild introduces the Aff-wild2 database as its cornerstone dataset, a compilation that encompasses substantial contributions from a series of studies [35, 16, 12, 17, 11, 18, 19, 13, 9, 10, 14, 15]. In the context of the EXPR Classification Challenge, the Aff-wild2 database, which is audiovisual in nature, comprises approximately 548 videos, culminating in around 2.7 million frames. These frames are meticulously annotated to reflect the six fundamental expressions—anger, disgust, fear, happiness, sadness, surprise—alongside a neutral state and an additional ’other’ category to denote expressions or affective states beyond the basic ones.
Furthermore, our research incorporates data from external facial expression databases, notably AffectNet[24] and ExpW[40], to enrich our analysis. AffectNet, which aggregates around 1 million facial images sourced from the internet, assigns eleven distinct emotion and non-emotion categorical labels. For our purposes, we utilized images corresponding to the first seven categories. Meanwhile, the Expression in-the-Wild Database (ExpW) amasses 91,793 facial images procured via Google image search, with each image manually classified into one of the seven foundational expression categories.
In this study, we achieved a balanced dataset by uniformly extracting 8,000 labeled images for each category from the Aff-Wild2 collection. To further enhance the diversity of our dataset and bolster the generalizability of our model, we also randomly selected 8,000 images for each category from a merged dataset comprising AffecNet and ExpW. For images falling into the ’other’ category, we relied exclusively on the Aff-Wild2 dataset, given the absence of such images in the alternate databases. The compilation is rounded off with the inclusion of the remaining images from these datasets as unlabeled samples, serving to augment our dataset’s comprehensiveness.
FR Datasets.
In our work on Face Recognition Datasets, we employ the MS1MV2 dataset[8] as the source of unlabeled data. This dataset, a semi-automatically refined iteration of the MS-Celeb-1M dataset[7], was developed by ArcFace[5] and contains approximately 85,000 identities and 5.8 million images. For the purposes of our experiments, we utilized a subset of the MS1MV2 dataset, curated from InsightFace[6], by uniformly selecting one-third of the images. This resulted in a comprehensive subset encompassing 1.94 million images.
4.2 Setup
[…]
4.3 Metrics
In line with the competition requirements, we employ the average F1 score as our evaluation metric, which is robust against class frequency variations and particularly suitable for imbalanced class distributions. The calculation of the average F1 score is as follows:
$F_{1}^{c}=\frac{2\times Precision\times Recall}{Precision+Recall}$ (7)
$F1=\frac{1}{N}\sum_{c=1}^{N}F_{1}^{c}$ (8)
where $N$ represents the number of classes and $c$ means $c$-th class.
4.4 Results
4.5 Ablation Study
To validate the efficacy of our approach, we executed ablation studies on each component and strategy within our method, with the outcomes presented in Table.LABEL:table1 It is evident that the application of the SSL (Semi-Supervised Learning) technique significantly enhances the recognition performance by 13.31%. Further augmentation of the facial expression dataset can lead to an increase in accuracy. Moreover, incorporating the temporal encoder results in an additional 2.2% improvement in accuracy, underscoring the importance of temporal learning. Ultimately, through post-processing, the model achieves an accuracy rate of 45.43%.
5 Conclusion